An electron passes through a point 2.67 cm from a long straight wire as it moves at 30.1% of the speed of light perpendicularly toward the wire. The wire carries a current of 17.9 A. Find the magnitude of the electron's acceleration at that point.
Answer #1

Given,

r = 2.67 cm ; v = 30.1% c ; I = 17.9 A

We know that

c = 3 x 10^8 m/s

30.1% of c = 0.301 x 3 x 10^8 m/s = 9.03 x 10^7 m/s

We know that the magnetic force on the electron is:

F = q v B

B = u0 I/2 pi r

B = 4 pi x 10^-7 x 17.9/(2 pi x 0.0267) = 1.34 x 10^-4 T

F = 1.6 x 10^-19 x 9.03 x 10^7 x 1.34 x 10^-4 = 1.94 x 10^-15 N

F = ma => a = F/m

a = 1.94 x 10^-15/(9.1 x 10^-31) = 2.13 x 10^15 m/s^2

Hence, a = 2.13 x 10^15 m/s^2